# Bayesian Quantile Regression with Subset Selection: A Posterior Summarization Perspective

### v3

@@ Abstract
Quantile regression is a powerful tool in epidemiological studies where interest lies in inferring how different exposures affect specific percentiles of the distribution of a health or life outcome. Existing methods either estimate conditional quantiles separately for each quantile of interest or estimate the entire conditional distribution using semi- or non-parametric models. The former often produce inadequate models for real data and do not share information across quantiles, while the latter are characterized by complex and constrained models that can be difficult to interpret and computationally inefficient. Further, neither approach is well-suited for quantile-specific subset selection. Instead, we pose the fundamental problems of linear quantile estimation, uncertainty quantification, and subset selection from a Bayesian decision analysis perspective. For any Bayesian regression model, we derive optimal and interpretable linear estimates and uncertainty quantification for each model-based conditional quantile. Our approach introduces a quantile-focused squared error loss, which enables efficient, closed-form computing and maintains a close relationship with Wasserstein-based density estimation. In an extensive simulation study, our methods demonstrate substantial gains in quantile estimation accuracy, variable selection, and inference over frequentist and Bayesian competitors. We use these tools to identify and quantify the heterogeneous impacts of multiple social stressors and environmental exposures on educational outcomes across the full spectrum of low-, medium-, and high-achieving students in North Carolina.
Variable selection, decision theory, interpretable machine learning, Bayesian inference
@@ 1 Introduction
(see Section 2.4 for the nonlinear case). Estimated across $\tau$, the coefficients $\boldsymbol{\beta}(\tau)$ summarize how the covariates affect not only the location, but also the shape of the response distribution $Y\mid\boldsymbol{x}$. The unique insight from linear quantile regression comes from identifying predictors with heterogeneous effects $\beta_{j}(\tau)\neq\beta_{j}(\tau^{\prime})$ for some $\tau\neq\tau^{\prime}$. For instance, when analyzing education data (see Sections 1.1 and 5), it is important not only to identify the factors that impact educational outcomes, but also to determine whether those effects are different for low-, medium, or high-achieving students. Such heterogeneous, quantile-specific effects can have far-reaching implications for policy interventions. In general, quantile regression provides a robust and comprehensive view of the relationship between covariates and the response variable, leading to many important applications: medicine (Kottas and Gelfand, 2001), finance (Bassett and Chen, 2002), and environmental studies (Pandey and Nguyen, 1999), among many others.
@@
We first introduce our motivating application, and then review existing Bayesian and frequentist methods for quantile regression–both to showcase the successes in this area and to highlight the need for methodological advances, especially for our data analysis.
1.1 Social Stressors, Environmental Exposures, and Childhood Educational Outcomes
Childhood educational outcomes are affected by social stressors (e.g., poverty, structural racism, high unemployment) and environmental exposures (e.g., poor air quality, lead exposure), which often cumulate in the same communities (Miranda et al., 2007; Bravo et al., 2024). Thus, there is urgency to quantify and characterize these effects and initiate well-informed policy interventions, especially given the strong links between educational attainment and adult health. However, far less is known about the differential impacts of social stressors and environmental exposures on low-, medium-, and high-achieving students. Such information is critical for understanding the factors that shape educational outcomes and for recommending policy interventions, especially to support at-risk students.
[…]
There are several important aspects and challenges in this analysis. First, EoG test scores are not merely proxies for educational achievement; more directly, they help determine whether students are promoted to the next grade, participate in programming for intellectually gifted students, or receive access to specialized support and resources. Clearly, it is important to differentiate among the factors that impact low-, medium-, and high-achieving students. Previous analyses of these data have exclusively used (homoscedastic) mean regression for estimation and variable selection (Kowal et al., 2021; Kowal, 2022; Bravo et al., 2022), which is not capable of estimating heterogeneous, quantile-specific effects. Second, the collection of potential variables (Table 1) is moderately large with substantial correlations among the variables (see the supplementary material, Section F.1), since many of the social and environmental exposures appear concurrently. Consequently, subset selection is especially useful, but must be i) quantile-specific, ii) capable of handling correlations among variables, and iii) scalable in $n$ and $p$. Finally, because of the significant policy implications, it is imperative that our numerical summaries are interpretable, robust, and accompanied by full uncertainty quantification. Because our quantile regression analysis will enumerate estimates (effect directions and magnitudes), uncertainties, and selected subsets across many quantiles, this requires stability (or smoothness) and probabilistic coherency across quantiles.
Birth information
mEdu
Mother’s education group at the time of birth
(No high school diploma, High school diploma,
College diploma)
mRace
Mother’s race/ethnicity group (Non-Hispanic (NH) White,
NH Black)
BWTpct Birthweight percentile
mAge Mother’s age at the time of birth
Male Male infant? (1 = Yes)
Smoker Mother smoked? (1 = Yes)
NotMarried Not married at time of birth (1 = Yes)
NOPNC
Mother received pre-natal care before birth?
(1 = No prenatal care)
Weeks_Gestation Gestational period (in weeks)
Education/End-of-grade (EoG) test information
Reading_Score
Standardized score for the (chronologically first) 4th EoG
reading test
Blood lead surveillance
Blood_lead Blood lead level (micrograms per deciliter)
Social/Economic status
EconDisadvantage
Participation in Child Nutrition Lunch Program? (1 = Yes)
NDI Neighborhood Deprivation Index, at time of EoG test
RI Racial residential isolation, at time of EoG test
@@ 1.2 Separate Quantile Regressions
In general, Bayesian quantile regression with the AL likelihood faces significant limitations. First, there is no information-sharing among nearby quantiles, which results in excessively large posterior uncertainties and underpowered inference for $\boldsymbol{\beta}(\tau)$, especially for extreme quantiles near zero or one (see Sections 4-5). Second, an AL likelihood must be specified for each $\tau$, which induces distinct Bayesian models for the same data, typically with no attempt to reconcile or combine them. Finally, the AL likelihood often produces a substantially inadequate model for real data across many, if not all $\tau$, which undermines the interpretability of the resulting inferences. Kowal and Wu (forthcoming) proposed a transformation-based generalization of (3) to improve model adequacy, but did not address the previous two concerns.
@@ 1.4 Subset Selection in Quantile Regression
Among separate quantile regression methods, modifications to (2) have been developed for quantile-specific variable selection. Sparse estimates are obtained by appending a penalty term to (2), such as an $\ell_{1}$-penalty (Wu and Liu, 2009; Belloni and Chernozhukov, 2011; Wang et al., 2012; Lee et al., 2014). The sparsity among penalized quantile regression estimates is controlled by a tuning parameter, typically selected via cross-validation, which can be computationally intensive for the (penalized) objective (2). These methods do not provide uncertainty quantification and are limited in that common sparsity penalties i) introduce overshrinkage of nonzero effects and ii) severely restrict the subset search path. By design, these methods focus on selecting a single “best” subset. However, even for moderate $p$ with correlated covariates or weak signals, there are often many subsets that offer similar predictive accuracy. Thus, it can be misleading to report only a single “best” subset; consideration of a broader collection of “near-optimal” subsets can be more comprehensive and informative, but requires a robust subset search (Kowal, 2022).
@@
1.5 Quantile Regression Through Posterior Summarization
To address the gaps in quantile regression methodology, we develop a Bayesian decision analysis procedure for estimation, uncertainty quantification, and subset selection for linear quantile regression. Our methodology centers around first building a Bayesian model to capture salient features in the data, and then extracting targeted summaries of the model-based conditional quantiles. The general procedure is outlined in Algorithm 1.
Algorithm 1 Bayesian quantile regression, inference and subset selection using posterior summarization
@@ 1.5 Quantile Regression Through Posterior Summarization
    Fit a Bayesian regression model $\mathcal{M}$
    • Description: Fit a Bayesian model to $\{\boldsymbol{x}_{i},y_{i}\}_{i=1}^{n}$ and extract posterior samples of model-based conditional quantile functions at each $\boldsymbol{x}_{i}$ and any quantile $\tau$ of interest;
  2. 2a.
    Quantile estimation and uncertainty quantification (Sections 2.1-2.2)
    • Description: For any $\tau$ and subset $S\subseteq\{1,\dots,p\}$ of predictors, apply decision analysis to obtain quantile-specific linear coefficient estimates and uncertainty quantification;
  3. 2b.
    Subset search, filtration, and selection (Section 3)
    • Description: For any $\tau$, conduct a quantile-specific subset search, accumulate a family of subsets with strong predictive power, and summarize this family via i) a single subset that balances parsimony and predictive power and ii) measures of variable importance across all subsets in the family.
Our formulation introduces several overarching advances to Bayesian quantile regression. First, the machinery is compatible with any Bayesian regression model $\mathcal{M}$ for $Y\in\mathbb{R}$, including non-linear and non-parametric models (Pratola et al., 2020), Bayesian model averaging (Raftery et al., 1997), and Bayesian model stacking (Yao et al., 2018), among many others. Our framework remains compatible with separate Bayesian AL models and Bayesian simultaneous quantile regression models. However, the enhanced generality of our approach allows the Bayesian modeler to build $\mathcal{M}$ to be suitable for the observed data, rather than to satisfy certain quantile-specific modeling requirements, such as AL likelihoods or unwieldy constraints.
Next, we design a decision analysis that extracts point estimates of quantile-specific linear coefficients from the model $\mathcal{M}$ conditional quantile functions. Our approach features a quantile-focused squared error loss, which is motivated through its close connection with the Wasserstein distance between measures (Fréchet, 1948) and enables efficient, closed-form computation of optimal (in a decision analysis sense) linear coefficients for any quantile and any subset of predictors. Crucially, the model $\mathcal{M}$ conveys regularization, both in the traditional sense (i.e., shrinkage of extraneous coefficients to zero) and via smoothness and information-sharing across nearby quantiles. Thus, while $\mathcal{M}$ need not feature linear quantiles (1), the conditional quantiles under a singular model $\mathcal{M}$ are probabilistically coherent and typically smooth across $\tau$.
@@ 2 Bayesian Decision Analysis for Quantile Regression
Given paired data $\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{n}$, consider a Bayesian regression model $\mathcal{M}$ for response variable $Y\in\mathbb{R}$ parameterized by $\boldsymbol{\theta}$. Suppose $\mathcal{M}$ has been curated to best capture the conditional distribution of $Y\mid\boldsymbol{x}$, which may include nonlinearity, skewness, and heteroscedasticity, among many other features.
@@
Given a prior $p(\boldsymbol{\theta})$, Bayesian inference for (4) targets the posterior distribution $p(\boldsymbol{\theta}\mid\boldsymbol{y})$. Then, for any $\tau$, the model $\mathcal{M}$ posterior propagates uncertainty to the conditional quantiles via (5). Thus, we view $Q_{\tau}(Y\mid\boldsymbol{x},\boldsymbol{\theta})$ as a posterior functional. When the error quantile function $F^{-1}$ is smooth in $\tau$, then the implied model-based quantiles (5) inherit this smoothness. Assuming $\mathcal{M}$ is a valid probability model for $Y\mid\boldsymbol{x}$, $Q_{\tau}(Y\mid\boldsymbol{x}_{i},\boldsymbol{\theta})$ is guaranteed to avoid quantile crossing.
Leveraging the Bayesian model $\mathcal{M}$, our goal is to provide linear quantile estimates, uncertainty quantification, and subset selection. We design a decision analysis for each task. Central to this approach is a loss function $\mathcal{L}\{Q_{\tau}(Y\mid\boldsymbol{x},\boldsymbol{\theta}),\boldsymbol{x}^{\intercal}\boldsymbol{\delta}_{S}(\tau)\}$ that evaluates the accuracy of a quantile-specific linear action, $\boldsymbol{x}^{\intercal}\boldsymbol{\delta}_{S}(\tau)$, with active (i.e., nonzero) coefficients for a given subset $S\subseteq\{1,\ldots,p\}$ of covariates. Since the model-based quantiles $Q_{\tau}(Y\mid\boldsymbol{x},\boldsymbol{\theta})$ are unknown but inherit a posterior distribution under $\mathcal{M}$ through $p(\boldsymbol{\theta}\mid\boldsymbol{y})$, the optimal coefficients are obtained by minimizing the posterior expected loss:
@@
Thus, (6) provides linear quantile estimation, specific to each quantile $\tau$ and each subset $S$, under the model $\mathcal{M}$ and the loss $\mathcal{L}$. We note that the proposed decision analysis framework is not strictly limited to linear actions $\boldsymbol{x}^{\intercal}\boldsymbol{\delta}_{S}(\tau)$: it is possible to extract quantile-specific trees or additive functions of $\boldsymbol{x}$ (see Section 2.4).
A principal task is to specify the loss function $\mathcal{L}$. We propose a quantile-focused squared error loss, aggregated over the covariate values $\{\boldsymbol{x}_{i}\}_{i=1}^{n}$:
@@
The loss function (7) is exceptionally convenient for point estimation (Section 2.1), uncertainty quantification (Section 2.2), and subset selection (Section 3), and enables closed-form estimation and efficient subset search strategies. More formally, we establish deeper theoretical justification for (7) via connections to Wasserstein-based density regression (Section 2.3).

### v4

# Bayesian Quantile Regression with Subset Selection: A Decision Analysis Perspective

@@ Abstract
Quantile regression is a powerful tool for inferring how covariates affect specific percentiles of the response distribution. Existing methods either estimate conditional quantiles separately for each quantile of interest or estimate the entire conditional distribution using semi- or non-parametric models. The former often produce inadequate models for real data and do not share information across quantiles, while the latter are characterized by complex and constrained models that can be difficult to interpret and computationally inefficient. Neither approach is well-suited for quantile-specific subset selection. Instead, we pose the fundamental problems of linear quantile estimation, uncertainty quantification, and subset selection from a Bayesian decision analysis perspective. For any Bayesian regression model—including, but not limited to existing Bayesian quantile regression models—we derive optimal point estimates, interpretable uncertainty quantification, and scalable subset selection techniques for all model-based conditional quantiles. Our approach introduces a quantile-focused squared error loss that enables efficient, closed-form computing and maintains a close relationship with Wasserstein-based density estimation. In an extensive simulation study, our methods demonstrate substantial gains in quantile estimation accuracy, inference, and variable selection over frequentist and Bayesian competitors. We use these tools to identify and quantify the heterogeneous impacts of multiple social stressors and environmental exposures on educational outcomes across the full spectrum of low-, medium-, and high-achieving students in North Carolina.
Variable selection, interpretable machine learning, Bayesian inference, robust regression
@@ 1 Introduction
(a nonlinear version is in Section 2.4). Estimated across $\tau$, the coefficients $\boldsymbol{\beta}(\tau)$ summarize how the covariates affect not only the location, but also the shape of the response distribution $Y\mid\boldsymbol{x}$. The unique insight from linear quantile regression comes from identifying predictors with heterogeneous effects $\beta_{j}(\tau)\neq\beta_{j}(\tau^{\prime})$ for some $\tau\neq\tau^{\prime}$. This capability is essential when the covariates affect higher order moments or other distributional features of $Y\mid\boldsymbol{x}$. For instance, when analyzing education data (see Section 5), it is important not only to identify the factors that impact educational outcomes, but also to determine whether those effects are different for low-, medium, or high-achieving students. Such heterogeneous, quantile-specific effects can have far-reaching implications for policy interventions. In general, quantile regression provides a robust and comprehensive view of the relationship between covariates and the response variable, leading to many important applications: medicine (Kottas and Gelfand, 2001), finance (Bassett and Chen, 2002), and environmental studies (Pandey and Nguyen, 1999), among many others.
@@
The confluence of these demands creates a challenging environment for general statistical procedures. Existing approaches diverge into separate paradigms, each with their own limitations (see Sections 1.1 and 1.2). With this in mind, we seek to
  1. 1.
    Develop a decision analysis framework for quantile regression that is foundational, coherent, and compatible with any Bayesian regression model.
Our decision analysis integrates quantile regression into a data-centric Bayesian workflow that prioritizes suitable modeling of observed data, rather than adherence to certain quantile modeling requirements (see Sections 1.1 and 1.2). This broadens the utility of quantile regression, since it can be incorporated alongside traditional posterior summaries such as posterior means and credible intervals for model parameters. At the same time, the foundational and general nature of the decision analysis provides a new path to
  1. 2.
    Rigorously unify divergent approaches for Bayesian quantile regression.
From this perspective, the decision analysis provides essential context for comparing and contrasting the large array of existing Bayesian quantile regression models. More specifically, the inferential goal of the decision analysis is to
  1. 3.
    Deliver optimal (in a decision theory sense) point estimates and uncertainty quantification for conditional quantile functions and quantile-specific linear coefficients.
This goal is shared by many quantile regression approaches, and thus presents an opportunity for competitive evaluations (Section 4). More uniquely, via decision analysis we
  1. 4.
    Provide efficient algorithms for quantile-specific subset selection.
Subset selection is a daunting task—especially for quantile regression. Our decision analysis leverages state-of-the-art search algorithms and subset selection strategies for mean regression, and neatly adapts them to quantile regression.
Before expanding upon these goals and contributions, we first review existing Bayesian and frequentist methods for quantile regression—both to showcase the successes in this area and to highlight the need for methodological advances.
@@ 1.1 Separate Quantile Regressions
In general, Bayesian quantile regression with the AL likelihood faces significant limitations. First, there is no information-sharing among nearby quantiles, which results in excessively large posterior uncertainties and underpowered inference for $\boldsymbol{\beta}(\tau)$, especially for extreme quantiles near zero or one (see Sections 4-5). Second, an AL likelihood must be specified for each $\tau$, which induces distinct Bayesian models for the same data, typically with no attempt to reconcile or combine them. Finally, the AL likelihood often produces a substantially inadequate model for real data across many, if not all $\tau$, which undermines the interpretability of the resulting inferences. Kowal and Wu (2024) proposed a transformation-based generalization of (3) to improve model adequacy, but did not address the previous two concerns.
@@ 1.3 Subset Selection in Quantile Regression
Among separate quantile regression methods, modifications to (2) have been developed for quantile-specific variable selection. Sparse estimates are obtained by appending a penalty term to (2), such as an $\ell_{1}$-penalty (Wu and Liu, 2009; Belloni and Chernozhukov, 2011; Wang et al., 2012; Lee et al., 2014). The sparsity among penalized quantile regression estimates is controlled by a tuning parameter, typically selected via cross-validation, which can be computationally intensive for the (penalized) objective (2). These methods do not provide uncertainty quantification and are limited in that common sparsity penalties i) introduce overshrinkage of nonzero effects and ii) severely restrict the subset search path. Critically, these methods focus on selecting a single “best” subset. However, even for moderate $p$ with correlated covariates or weak signals, there are often many subsets that offer similar predictive accuracy. Thus, it can be misleading to report only a single “best” subset; consideration of a broader collection of “near-optimal” subsets can be more comprehensive and informative, but requires a robust subset search (Kowal, 2022).
@@
1.4 Overview of the Proposed Approach
To address the gaps in quantile regression methodology, we develop a Bayesian decision analysis for estimation, uncertainty quantification, and subset selection for quantile regression. Our methodology centers around first building a Bayesian model to capture salient features in the data, and then extracting targeted summaries of the model-based conditional quantiles. The general procedure is outlined in Algorithm 1.
Algorithm 1 Bayesian decision analysis for quantile regression
@@ 1.4 Overview of the Proposed Approach
    Fit a Bayesian regression model $\mathcal{M}$ (e.g., (4))
    Inputs: paired data $\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{n}$
    Output: posterior distribution (e.g., posterior samples) of the model-based conditional quantiles $\{Q_{\tau}(Y\mid\boldsymbol{x}_{i},\boldsymbol{\theta})\}_{i=1}^{n}$ (e.g., (5)) for each quantile $\tau$
  2. 2.
    Conduct Bayesian decision analysis for quantile regression:
    1. (a)
      Compute quantile-specific point estimates (Sections 2.1 and 2.4)
      Output: optimal actions (e.g., (8)) that minimize the quantile-focused loss (7) for each quantile $\tau$ (nonlinear case: Section 2.4)
    2. (b)
      Provide quantile-specific posterior uncertainty quantification (Section 2.2)
      Output: distribution for the posterior action (9) for each quantile $\tau$
    3. (c)
      Search, filter, and select quantile-specific subsets (Section 3)
      Output: quantile-specific acceptable family (17) that collects near-optimal subsets; quantile-specific smallest acceptable subset (18); and quantile-specific variable importance metrics (19).
Our formulation introduces several overarching advances to Bayesian quantile regression. First, the machinery is compatible with any Bayesian regression model $\mathcal{M}$ for $Y\in\mathbb{R}$, including non-linear and non-parametric models (Pratola et al., 2020), Bayesian model averaging (Raftery et al., 1997), and Bayesian model stacking (Yao et al., 2018), among many others. Our framework remains compatible with separate Bayesian AL models and Bayesian simultaneous quantile regression models, thus offering a unified framework for these divergent approaches. However, the enhanced generality of our approach allows the Bayesian modeler to build $\mathcal{M}$ to be suitable for the observed data, rather than to satisfy certain quantile-specific modeling requirements, such as AL likelihoods or unwieldy constraints.
Next, we design a decision analysis that extracts point estimates of quantile-specific coefficients from the model $\mathcal{M}$ conditional quantile functions. Like any decision analysis, we must select a loss function; our approach features a quantile-focused squared error loss. This choice is motivated through its close connection with the Wasserstein distance between measures (Fréchet, 1948) and enables efficient, closed-form computation of optimal (in a decision analysis sense) coefficients for any quantile and any subset of predictors. Crucially, the model $\mathcal{M}$ conveys regularization, both in the traditional sense (i.e., shrinkage of extraneous coefficients to zero) and via smoothness and information-sharing across nearby quantiles. Thus, while $\mathcal{M}$ need not feature linear quantiles (1), the conditional quantiles under a singular model $\mathcal{M}$ are probabilistically coherent and typically smooth across $\tau$.
@@
The remainder of the paper is organized as follows. In Section 2, we introduce posterior decision analysis for quantile regression. Section 3 describes our quantile-specific subset search and selection procedure. We provide a simulation study in Section 4 to evaluate the proposed methodology for prediction, inference, selection, and quantile crossing. In Section 5 we conduct a quantile regression analysis using data on end-of-grade test scores for children in North Carolina. We conclude in Section 6. Supplementary materials include proofs to all results, Bayesian model specifications, additional simulation results, and an R package [^1] Codes implementing the proposed subset search and selection methodology are available at https://github.com/jfeldman396/QRSubsets implementing the proposed methodology.
@@ 2 Bayesian Decision Analysis for Quantile Regression
Given paired data $\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{n}$, consider a Bayesian regression model $\mathcal{M}$ for response variable $Y\in\mathbb{R}$ parameterized by $\boldsymbol{\theta}$. Consistent with a data-centric Bayesian workflow, suppose $\mathcal{M}$ has been curated to best capture the conditional distribution of $Y\mid\boldsymbol{x}$, which may include nonlinearity, skewness, and heteroscedasticity, among many other features.
@@
Given a prior $p(\boldsymbol{\theta})$, Bayesian inference for (4) targets the posterior distribution $p(\boldsymbol{\theta}\mid\boldsymbol{y})$. Then, for any $\tau$, the model $\mathcal{M}$ posterior propagates uncertainty to the conditional quantiles via (5). Accordingly, we view $Q_{\tau}(Y\mid\boldsymbol{x},\boldsymbol{\theta})$ as a posterior functional. When the error quantile function $F^{-1}$ is smooth in $\tau$, then the implied model-based quantiles (5) inherit this smoothness. Assuming $\mathcal{M}$ is a valid probability model for $Y\mid\boldsymbol{x}$, $Q_{\tau}(Y\mid\boldsymbol{x}_{i},\boldsymbol{\theta})$ is guaranteed to avoid quantile crossing.
Leveraging the Bayesian model $\mathcal{M}$, our goal is to provide linear quantile estimates, uncertainty quantification, and subset selection. We design a decision analysis for each task. Central to this approach is a loss function $\mathcal{L}\{Q_{\tau}(Y\mid\boldsymbol{x},\boldsymbol{\theta}),\boldsymbol{x}^{\intercal}\boldsymbol{\delta}_{S}(\tau)\}$ that evaluates the accuracy of a quantile-specific linear action, $\boldsymbol{x}^{\intercal}\boldsymbol{\delta}_{S}(\tau)$, with active (i.e., nonzero) coefficients for a given subset $S\subseteq\{1,\ldots,p\}$ of covariates. Nonlinear actions such as quantile-specific trees or additive functions are also compatible within this framework (see Section 2.4).
Since the model-based quantiles $Q_{\tau}(Y\mid\boldsymbol{x},\boldsymbol{\theta})$ are unknown but inherit a posterior distribution under $\mathcal{M}$ through $p(\boldsymbol{\theta}\mid\boldsymbol{y})$, the optimal coefficients are obtained by minimizing the posterior expected loss:
@@
Thus, (6) provides linear quantile estimation, specific to each quantile $\tau$ and each subset $S$, under the model $\mathcal{M}$ and the loss $\mathcal{L}$.
These steps—fitting a Bayesian model $\mathcal{M}$ and then minimizing a posterior expected loss—are the foundational and uncontroversial components of a Bayesian decision analysis. The principal task is to specify the loss function $\mathcal{L}$. We propose a quantile-focused squared error loss, aggregated over the covariate values $\{\boldsymbol{x}_{i}\}_{i=1}^{n}$:
@@
The loss function (7) is exceptionally convenient for point estimation (Section 2.1), uncertainty quantification (Section 2.2), and subset selection (Section 3), with closed-form estimation and efficient subset search strategies. More formally, we establish deeper theoretical justification for (7) via connections to Wasserstein-based density regression (Section 2.3).